OpenStudy (anonymous):

7 + x = 1/2 times (4x - 2) ((show work))

OpenStudy (shinalcantara):

cross multiplying 2 you'll have (7 + x) (2) = (4x - 2) distribute 2 7(2) + x(2) = 4x -2 14 + 2x = 4x - 2 transposing the values you'll have 2x - 4x = -2 - 14 -2x = -16 x = 8